OpenStudy (anonymous):

let f(x)=(x-2)/(3x+3) find an equation for the tangent line to the graph of f at x=10

OpenStudy (anonymous):

f(10)=8/33 f'(x)=1/(x+1)^2 f'(10)=1/121 y-y0=m(x-x0) thus equation is: y-8/33=(1/121)(x-10)